every boy had 5 hotdogs and 3 sodas. every girl had 3 hotdogs and 2 sodas. together, the children ate 74 hotdogs and 46 sodas. how many children were there?
Answer by stanbon(75887)

every boy had 5 hotdogs and 3 sodas. every girl had 3 hotdogs and 2 sodas. together, the children ate 74 hotdogs and 46 sodas. how many children were there?
---
Equations:
hotdog Eq: 5b + 3g = 74
soda Eq: 3b + 2g = 46
-----------------------
Multiply thru 1st by 2
Multiply thru 2nd by 3
----------------------------
10b + 6g = 148
9b + 6g = 138
-------------------
Subtract and solve for "b":
b = 10 (# of boys)
------
Solve for "g":
3b + 2g = 46
3*10 + 2g = 46
2g = 16
g = 8 (# of girls)
----
b + g = 10+8 = 18 (# of children)
==================================
Cheers,
Stan H.
============